LBTO Technical/D-Night

OSA: D.G.Huerta

ISp: J. Morris

Manager: J. Urban

ISA: B. Rothberg, D. Thompson (Tucson)

=====================================================

UT 01:00 - Thick clouds overhead, humidity 8%

UT 01:11 - Dome open - trying to get some flux on WFS, not looking promising

UT 02:37 - dome closed thick clouds

UT 04:00 - Why is there a Dalek wandering around on the 5th floor?

UT 06:06 - Going to run tests for standard calibrations to get itimes and settings

UT 06:12 - reinitializing as I'm getting lots of errors

UT 06:22 - the bottom half of slit seems illuminated - moving cal unit back out - this turned out to be the AGW on axis - I keep forgetting this happens!

UT 06:26 - initializing the calibration unit

UT 06:40 - possible break, we are opening up now

===============================================================================

UT 06:51 - HD 200527 - H-band Library star

UT 06:59 - GUIs froze, killed java and reopemed

UT 07:00 - got error TASK:failed to center slit positions, Message: failed to select/find requested Aladin Plane: S_slit_trg1 Exception Null

UT 07:01 - taking another thru slit

UT 07:02 - freezing meant 3 images were taken and I didn't know - it jumped to 18

UT 07:04 - spectroscopic exposures being taken - conditions might not be that bad - 0020-0023 are data

UT 07:12 - moving to Telluric

UT 07:18 - appears to be two stars, one that landed in slit i'm assuming is it

UT 07:19 - taking data

======================================================

UT 07:25 - Moving to M.Wagner M31 target

UT 07:28 - Can't find guidestar- checking pointing

UT 07:31 - pointing updated - back to field seeing 0.8"

-first image is 0031

UT 07:37 - seeing median 0.6"

UT 07:42 - Java has frozen up again

UT 07:56 - Moving to sky field for M31 M. Wagner target

UT 08:02 - median seeing 0.5-0.6" in guider

UT 08:21 - median seeing 0.6 - last image in set is 0060

========================================

UT 08:23 - need to unwrap Azimuth - going to Rothberg LIRG UGC 6

UT 08:27 - slewing to UGC 6

UT 08:29 - had to put mask in FPU during unwrap - LUCI2 got confused again - kept MASK in FPU even though not suppsoed to - 0061,0062 useless

UT 08:32 - initialized all

UT 08:33 - resending preset

UT 08:43 - acq complete science exposure started

UT 08:51 - on/off "not stable" on the grating voltage

UT 09:45 - last exposure is airmass 2.2 seeing 1.6 in guider

UT 09:49 - Moving to Telluric for UGC 6

UT 09:55 - Telluric is a double! I think I have it centered on the right one, but I'm going to acquire again and take spectra of both to be safe

UT 10:00 - Resending same script to take spectra of other star - centered at y=990 to avoid persistance from pervious observation

UT 10:06 - taking spectra of 2nd star

=========================================

UT 10:11 - INAF Antoniucci v1118_Orionis Spectroscopy

UT 10:14 - guider says 0.7" seeing - first exposure 0083

UT 10:15 - PI selected 0.75" slit which we don't have - consulting Felice

UT 10:19 - replaced 0.75" with 1.00" slit

UT 10:28 - taking data

UT 10:34 - median seeing guider 0.7"

UT 10:46 - median seeing guide 0.7" - clouds rolling in, currently pointed between clouds

UT 10:54 - executing imaging script v1118_orio_img.xml

UT 10:56 - pointing check

UT 11:01 - pointing updated - running img script

UT 11:05 - clicked science instead of telluric, skip 101,102

UT 11:08 - now moving to Telluric field HIP 337323

UT 11:15 - taking spectra telluric

UT 11:21 - Moving to SDSS_J0510 (modified to 1" slitwidth) - shares telluric with HIP 337323

UT 11:22 - Seeing 0.6 median in guider

UT 11:27 - science exposures started

UT 11:37 - clouds really rolling in -good thing target is bright

UT 11:48 - guider showing increased background flux from clouds and scattered light of Moon

UT 11:59 - SDSS_J0510_img imaging script

UT 12:02 - guiding is barely hanging on

UT 12:04 - 0125 guiding is gone - no star found

========================================

UT 12:09 - closing DOME - prep for calibrations

UT 12:22 - moving Cal unit in

UT 12:24 - error in script - mask was never put into place

UT 12:29 - running INAF script 1" slit with zJHK 1.175 and 1.93 CWL

UT 12:34 - again, script and LUCI-2 didn't match - everything up to 145 is junk.

UT 12:35 - 146 - start of 1" INAF zJHK 200 1.175 and 1.93 CWL cals

UT 12:41 - zJ 3sec for 1.175 CWL 1" G200 counts are low - going to re-run just this portion with exptime =6sec

UT 12:44 - GUI shows almost no info during cals (146-197)

UT 12:59 - Running 210 K 0.5" slit 2.2 CWL 0198-

UT 01:08 - skipped over the lights off went to lights on for flats - counts are low 2k for 3.0 sec 198-218 are useless

UT 01:19 - testing exposure times rather than running scripts

UT 01:28 - shutting down LUCI-2 now, 3.75, blind filters and no mask in FPU, mirror

UT 01:34 - DONE

-- %USERSIG{BarryRothberg - 2015-10-28}%

Comments

 
Topic revision: r6 - 28 Oct 2015, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback